Tell how many solutions the system has. Do not actually solve.
2x - y = 5
-4x + 2y = -18

Found 2 solutions by fractalier, rajagopalan:
Answer by fractalier(6550)

The system
2x - y = 5
-4x + 2y = -18
has no solutions.
These lines are parallel.
The system is called inconsistent.

Answer by rajagopalan(174)

2x - y = 5
-4x + 2y = -18
LHS is the same -4x+2y =-2(2x-y)
So both LHS are same. This can not construe a system of equations.
If you consider each eqn individually u can have infinite solutions.
Good luck